OpenStudy (anonymous):

3(x+5)-10=-2(x+10). what is the solution?

OpenStudy (anonymous):

First, multiply out both sides 3(x+5)-10=-2(x+10) 3x + 15 - 10 = -2x - 20 Then simplify where possible 3x + 5 = -2x - 20 Then move all the x to one side 5x = -25 Finally simplify to get just x x = -5

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem